OpenStudy (anonymous):

The hieght of a rocket above the ground is modelled by the quadratic function h(t) = -4t^2 + 32t, where h(t) is the height in metres, t seconds after the rocket was launched. a. How long will the rocket be in the air? How do you know? b. How high will the rocket be after 3s?

OpenStudy (anonymous):

Solve for the zeroes of h(t), and from those you should have the time it lands. Plugging in 3 to h(3) will give you the height after 3 seconds.

OpenStudy (anonymous):

Well I don't know if my procedure is right but the answer is correct: a. h(t) = -4t^2 + 32t 0 = -4t^2 + 32t -32t = -4t^2 ----- ------ -4t -4t 8 = t

OpenStudy (anonymous):

So to solve the zeros I use the factored form?

OpenStudy (anonymous):

Yes, or the quadratic equation.

OpenStudy (anonymous):

How do i use the factored form when it's f(x) = ax^2 + bx?

OpenStudy (anonymous):

x = (-b +- sqrt(b^2 - 4 * a * c)) / (2 * a)

OpenStudy (anonymous):

Plug in the coefficients and you should have the time(s) it is at ground level.

OpenStudy (anonymous):

ohh, so I use the quadratic formula? But how about if I want to use the factored form?

OpenStudy (anonymous):

-4t(t-8) is factored.

OpenStudy (anonymous):

also to get c in the quadratic formula I must c= -b/2a? or c = (-b/2)^2

OpenStudy (anonymous):

-4t(t-8) = 0 => t = 0, 8.

OpenStudy (anonymous):

ohhhh.